An experiment has four outcomes O1,O2,O3, and O4 with corresponding weights w1,w2,w3, and w4 where w1=7/k, w2=7/k,w3=6/k, and w4=1/k. What is the value of k? What is Pr({O1,O4})?
Answer #1

The sum of all the probabilities should be 1. Hence,

w1 + w2 + w3 + w4 = 1

7/k + 7/k + 6/k + 1/k = 1

21/k = 1

k = 21

Pr({O1, O4}) = w1 + w4 = 7/21 + 1/21 = 8/21